# Pauli Stabilizer Models for Gapped Boundaries of Twisted Quantum Doubles and Applications to Composite Dimensional Codes

### v3

@@ V.1 Examples
Figure 35: Minimal stabilizer model for the $\mathbb{Z}_{2}$ on a cylinder with two rough boundaries. The double arrows indicate boundary conditions where top edges are identified with bottom edges. The semi-transparent diamonds are identified with the ones in the bottom and should not be counted. Thus, there are $15$ edges, 99 plaquettes, and $6$ stars in this figure.
Let us track the changes from Example V.2. The edges went from $18$ to $15$ and thus $r_{D}=2^{18-15}=2^{-3}$. There are only order $2$ operators. The number of plaquettes did not change while the number of stars decreased by $2$, so $\Delta n_{2}=-2$. Finally, while the dependencies were $l_{2}=0$ in the surface code, here the product of all plaquettes gives us the identity $\prod_{p}B(p)=\mathbb{I}$, so we have $l^{\prime}_{2}=1$. This gives $\Delta l_{2}=1$. We then compute the GSD from Eq. (59) as:
$GSD^{\prime}=\frac{2^{-3}}{2^{-2-(-1)}}\operatorname{GSD}=1$ (61)
Example V.4 ($\mathbb{Z}_{2}$ on a Torus through folding a cylinder).
@@
This method of constructing a minimal model to count the degeneracy becomes especially useful for more complicated configurations. All one needs to do is to make sure no product of the operators gives identity. This is the same condition that all the operators are independent. An interesting domain wall which was discussed is the $e\leftrightarrow m$ of the toric code IV.4. We can compute the GSD for this model using this simple model construction.
Example V.5 ($\mathbb{Z}_{2}$ on a Torus with a contractible $e\leftrightarrow m$ twist).
A minimal example for the toric code with a contractible twist is depicted in Fig. 37. We perform Alg. 1 from the toric code and carefully count what changes. We remove 22 AA stabilizers and 22 $B$ stabilizers and add one $T$, $U$ and $Q$. They are all of order $m=2$. In total, we have $\Delta n_{2}=3-4=-1$.
@@
This is correct because the codespace is the same as before, although a global phase factor 1-1
may now be applied due to the presence of the lone twist [53].
@@
Example V.6 ($\mathbb{Z}_{2}$ on a torus with $N$ twists).
@@
Example V.7 ($\mathbb{Z}_{2}$ on a Torus with a non-contractible twist).
@@
This example follows the result of Ref. 83 where it was shown that adding a non-contractible twist gauges the toric code’s $\widetilde{X_{1}}\widetilde{Z_{2}}$ degree of freedom, hence adding one logical constraint resulting in a single logical qubit [83].
Remark V.1.
@@
Having developed Alg. 1 to gauge stabilizer codes, and Eq. 59 to count and update degeneracies, we are now in a position to prove the results presented in Sec. III.3, which concern twisted quantum doubles coexisting. Since the twisted quantum doubles can be formed by condensation from quantum doubles [40], the degeneracy of such systems can be proved with minimal illustrative models. Continuing with the arguments above, we begin by considering a $D(\mathbb{Z}_{4})$ toric code with $GSD=4^{2}=16$. We now proceed to condense finite spatial regions and determine the GSDs.
Example V.8 ($\mathbb{Z}_{4}$ with a contractible $DS$ patch).
Start from the $D(\mathbb{Z}_{4})$ toric code and condense just one region that contains the DS phase. Any model with a single contractible (meaning it does not wrap around either handle of the torus) patch will have the same degeneracy as this construction. Before condensing the DS patch, we had $\prod_{v}A_{v}=1$ and $\prod_{p}B_{p}=1$, which meant that one of the plaquettes and one of the star operators is dependent for a total of two order four dependencies ($l_{4}=2$) and zero order two dependencies ($l_{2}=0$) such that $l=l_{2}+l_{4}=2$. To condense a DS patch, we need to add two $X^{2}Z^{2}$ ribbons, as in Fig. 40. Afterwards, we remove three-star operators and three plaquette operators that do not commute with the $X^{2}Z^{2}$ ribbons. We then add products of the removed operators that commute with the new stabilizers. These will be the $F(v,p)$ and $B^{2}(p)$ operators Eq. (22). We will add three $F$ operators on the removed vertices and three $B^{2}(p)$ operators on the removed plaquettes.
@@
Note that we have again two new constraint relations. First, we have $\prod_{v,p}A(v)B(p)F(v,p)=1$ because this is the same as multiplying all star and plaquette operators before adding the DS patch. Second, we have $\left(\prod_{p\in DS}B^{2}(p)\right)\left(\prod_{p\notin DS}B(p)\right)^{2}=1$. Here, the product is over the new plaquettes, which live inside or adjacent to the right or top of the DS patch, with the square of the rest of the old plaquettes that were not removed. Thus, $l^{\prime}=2$ and $r_{l}=2/2=1$. The crucial observation is that the first constraint will remove an operator with order four, for example, one $F(v,p)$, while the second constraint means we are double-counting an operator of order $2$, taken here as $B^{2}(p)$.
@@
Example V.9 ($\mathbb{Z}_{4}$ on a Torus with NN DSDS patches).
Starting with example V.8 with one DSDS patch, we can add one more patch as shown in Fig. 41.
@@
Table 5: table summarizing changes from Z4 code to Z4 and DS half and half contractible code.
@@
Example V.10 (Torus with half $\mathbb{Z}_{4}$ and half $DS$ phase).
Let us start with a $\mathbb{Z}_{4}$ on a torus. We want to calculate the degeneracy of a phase where half of the torus is a $\mathbb{Z}_{4}$ QD and the other half is in the Doubled Semion (DS) phase. Since here the DS lives on a cylinder, the boundary conditions will be different from just a DS patch.
@@
(b) Minimal lattice model for the configuration on the right. Compare with Example VI.3, which uses a different method.
@@
This can be understood intuitively since the full $\mathbb{Z}_{4}$ loops can no longer traverse the torus as some of them get identified or confined when they pass through the DS phase. Another way to calculate the $\operatorname{GSD}$ will make this clearer in VI. We saw in example V.8 that when the DS phase occupied a contractible region, the degeneracy was the same as without the DS phase. The stabilizer counting using minimal models is then a powerful tool for exploring modifications to topological orders in a sequential way. Paired with Alg. 1, one can imagine using a computer program to automate the construction of topological quantum error-correction codes.
VI Generic Code Construction
Dimensional counting and explicit evaluation of trace formulas are useful for deriving the logical properties, such as ground state degeneracy, from microscopic configurations. A complementary, top-down rather than bottom-up, and intuitive methodology, involving only high-level macroscopic information, is to count the number of non-contractible logical loops of anyons in our model. Remember that the ribbon operators are precisely the elementary operators that only have excitations at their two ends. Thus, if the two ends form a closed ribbon operator, or condense into a boundary or domain wall, then this operator commutes with the Hamiltonian, leaving it in its initial eigenspace. The operator may still act non-trivially on the ground state subspace, thus permuting ($\widetilde{X}$) or phasing ($\widetilde{Z}$) ground state sectors. On the other hand, if the closed loop is contractible, it will act trivially ($\widetilde{\mathds{1}}$). We then have to count how many independent nontrivial loops we have. This is best illustrated by an example.
Example VI.1 ($D(\mathbb{Z}_{2})$ on a Cylinder with two smooth boundaries).
Let’s take the cylinder with both smooth boundary conditions as shown in Fig. 43. We have two non-contractible loops around the cylinder, but one of them, namely the $X$ loop, will have a trivial action on the ground state as it can be condensed on either smooth boundaries (it is a product of stabilizers). This means that only the $Z$ loop around the cylinder is non-trivial. We also have an $X$ string between the two smooth boundaries of the cylinder. However, the $Z$ loop and the $X$ string anticommute, and we have $\operatorname{GSD}=2$. Evidently, all we needed is the information about the topological surface the theory lives on, along with the types of boundaries it has.
@@ VI Generic Code Construction
VI.1 Topological Degeneracy
In the mathematical field of topology, it is important to systematically classify how arbitrary surfaces are constructed in terms of simpler primitive structures. For example, how standard spheres and holes can be glued to form any connected and simply-connected 2-complex is a well-studied problem [85]. In this spirit, this section sketches how the loop counting argument can be generalized to the case of various topological orders supported on a $g$-genus 2D orientable manifold, which is allowed to have boundaries [64]. The construction relies on two facts. First, as shown in Fig. 45, any orientable 2D manifold can be decomposed into caps, cylinders, and pants. Second, there is a correspondence between domain walls (boundaries) and a tunneling or transfer matrix that describes the domain wall’s action on anyons.
@@ VI.1 Topological Degeneracy
Systematically, the method can be described as follows. First, we begin by taking as an input a 2D surface $\Sigma$ that can have multiple spatial boundaries $\partial\Sigma_{i}$. The surface may also be divided into different topological phases with appropriate domain walls $W^{j}$ between them. Boundaries are described as domain walls with the vacuum $W^{i}$ for uniformity. Second, we decompose the 2D surface using pants decomposition into the 2D components described in Fig. 45. This will isolate different bulk phases that have domain walls between them into different elementary 2D components. Multiple pants decompositions exist for the same configuration, and it is not crucial to use the minimal one. The different decompositions are related by a set of topological moves that will naturally not change the topology of the surface nor affect the $\operatorname{GSD}$ calculated here [86, 64]. We also do not discuss how this decomposition is formally carried over, since in all the examples, finding a decomposition is relatively easy. Third, we attach to every 2D component its appropriate tensor from Fig. 45. The tensors take values in the anyon labels of the bulk (boundary) of their respective topological phases. In this step, we also need to choose labels for the indices. Two indices that are connected should have the same label. In the end, no free indices should remain. Fourth, the $\operatorname{GSD}$ will be the scalar resulting from contracting the indices with the same label. This process is better illustrated with examples.
Example VI.2 ($\mathbb{Z}_{2}$ on a Torus with a non-contractible twist II).
@@
agreeing with Ref. 83 and our result using stabilizers model in Ex. V.7.
@@
Example VI.3 (Z4 with a non-contractible DS patch).
@@
which confirms the results of Ex. V.10.
@@
Example VI.4 ($D(\mathbb{Z}_{4})$ with a one non-contractible DS patch).
@@
All the domain walls between two $\mathbb{Z}_{4}$ phases are trivial and denoted here by $W^{1}$. The domain wall from $\mathbb{Z}_{4}$ to $DS$ is denoted by $W$, which is given in Ex. VI.3. Finally, we attach to the pants the $D(\mathbb{Z}_{4})$ fusion coefficients $N_{i,j}^{k}=\delta_{i+j,k}$ and to the cap the delta operator. The $\operatorname{GSD}$ is then:
@@
where in the second step we used $N_{i,a}^{a}=\delta_{i,1}$. In the last step, we just summed over the anyon types of $D(\mathbb{Z}_{4})$. This agrees with the results of Ex. V.8.
Example VI.5 ($D(\mathbb{Z}_{4})$ on a torus with a $N$ DS patches).
@@
We note that $W^{\dagger}_{x,1}$ is only non-zero when $x=1$ or $x=e^{2}m^{2}$ as can be seen from Eq. (75). We then have only two choices for the indices $b_{i}$, and there are N1N-1 of them (since $b_{N}$ is dependent). This gives $2^{N-1}$ choices. After picking the free N1N-1 b’s, picking any arbitrary choice for $a_{1}$ from the $16$ anyons of $\mathbb{Z}_{4}$ will determine the rest of $a^{\prime}s$ in addition to $b_{N}$. This gives $16\times 2^{N-1}$ choices. This agrees with Ex. V.9.
@@
Example VI.6 (DS Patches in a $\mathbb{Z}_{4}$ bulk with smooth external boundaries).
@@
In the second line, we enforced the delta functions. In the third line, since $a,c\in\{1,e^{2}m^{2}\}$, the only pure flux $b$ that can appear as (ac)(a-c) is the vacuum $1$. Finally, summing over the two choices of $a$ gives us $\operatorname{GSD}=2$. The case with $N$ patches can be treated similar to Example VI.5
Figure 50: Pants decomposition of the $\mathbb{Z}_{4}$ toric code with two contractible DS patches and an external smooth boundary.
@@ VII Conclusion
In this work, we have defined a new family of Abelian quantum error correcting codes. In addition to conventional boundaries of quantum doubles [62, 45], our codes contain finite-sized condensates of a twisted quantum double of the parent quantum double. We have explicitly illustrated this using the group $G=\mathbb{Z}_{4}$ and generalizing the DS global condensation[40] to a spatially local one. We derived the logical operators and showed that, in addition to whatever background logical operators of the parent $D(\mathbb{Z}_{4})$, N1N-1 logical qubits are added via $N$ patches. This result explicitly shows how topological qubits of distinct dimensions (two and four) can co-exist in a single code. We have described why, from the perspective of simplifying and improving error correction in the near term, this approach is desirable
@@
During the final stages of writing this manuscript, the authors became aware of Ref. [87] which analyzes phase transitions between DS and $\mathbb{Z}_{4}$ phases. While complementary, our work is focused on the quantum error-correcting codes derived by judicious use of these phases. It will be interesting to investigate the intersection of these two concepts in future studies.

### v4

@@ VI.1.1 Examples using Minimal Stabilizer Models
Figure 38: Minimal stabilizer model for the $\mathbb{Z}_{2}$ on a cylinder with two rough boundaries. The double arrows indicate periodic boundary conditions. Consequently, the semi-transparent diamonds should not be counted.
Let us track the changes from the layout in Fig. 37. The edges went from $18$ to $15$ and thus,
$r_{D}=2^{18-15}=2^{-3}.$
There are only order-$2$ operators. The number of plaquettes did not change while the number of stars decreased by $2$, so
$\Delta n_{2}=-2.$
Finally, while the dependencies were $l_{2}=0$ in the surface code, here the product of all plaquettes gives us the identity $\prod_{p}B(p)=\mathbb{I}$, so we have $l^{\prime}_{2}=1$. This gives
$\Delta l_{2}=1.$
We then compute the GSD from a direct application of Eq. (53) as:
$GSD^{\prime}=\frac{r_{D}}{2^{\Delta n_{2}-\Delta l_{2}}}=\frac{2^{-3}}{2^{-2-(1)}}\operatorname{GSD}=2$ (54)
Example VI.2 (Toric code through folding a cylinder).
@@
This method of constructing a minimal model to count the degeneracy becomes especially useful for more complicated configurations. All one needs to do is to make sure no product of the operators gives identity. This is the same condition that all the operators are independent. An interesting domain wall which was discussed is the $e\leftrightarrow m$ of the toric code V.4. We can compute the GSD for this model using this simple model construction.
Example VI.3 ($\mathbb{Z}_{2}$ on a Torus with a contractible $e\leftrightarrow m$ twist).
A minimal example for the toric code with a contractible twist is depicted in Fig. 40. We perform Alg. 2 from the toric code and carefully count what changes. We remove 22 AA stabilizers and 22 $B$ stabilizers and add one $T$, $U$ and $Q$. They are all of order $m=2$. In total, we have $\Delta n_{2}=3-4=-1$.
@@
This is correct because the codespace is the same as before, although a global phase factor 1-1 may now be applied due to the presence of the lone twist [21].
@@
Example VI.4 ($\mathbb{Z}_{2}$ on a torus with $N$ twists).
@@
Example VI.5 ($\mathbb{Z}_{2}$ on a Torus with a non-contractible twist).
@@
This example follows the result of Ref. 144 where it was shown that adding a non-contractible twist gauges the toric code’s $\widetilde{X_{1}}\widetilde{Z_{2}}$ degree of freedom, hence adding one logical constraint resulting in a single logical qubit [144].
Remark VI.1.
@@
Having developed Alg. 2 to gauge stabilizer codes, and Eq. 53 to count and update degeneracies, we are now in a position to prove the results presented in Sec. IV.4.4, which concern twisted quantum doubles coexisting. Since the twisted quantum doubles can be formed by condensation from quantum doubles [66], the degeneracy of such systems can be proved with minimal illustrative models. Continuing with the arguments above, we begin by considering a $D(\mathbb{Z}_{4})$ toric code with $GSD=4^{2}=16$. We now proceed to condense finite spatial regions and determine the GSDs.
Example VI.6 ($\mathbb{Z}_{4}$ with a contractible $DS$ patch).
Start from the $D(\mathbb{Z}_{4})$ toric code and condense just one region that contains the DS phase. Any model with a single contractible (meaning it does not wrap around either handle of the torus) patch will have the same degeneracy as this construction. Before condensing the DS patch, we had $\prod_{v}A_{v}=1$ and $\prod_{p}B_{p}=1$, which meant that one of the plaquettes and one of the star operators is dependent for a total of two order four dependencies ($l_{4}=2$) and zero order two dependencies ($l_{2}=0$) such that $l=l_{2}+l_{4}=2$. To condense a DS patch, we need to add two $X^{2}Z^{2}$ ribbons, as in Fig. 43. Afterward, we remove three-star operators and three plaquette operators that do not commute with the $X^{2}Z^{2}$ ribbons. We then add products of the removed operators that commute with the new stabilizers. These will be the $F(v,p)$ and $B^{2}(p)$ operators Eq. (19). We will add three $F$ operators on the removed vertices and three $B^{2}(p)$ operators on the removed plaquettes.
@@
Note that we have two new constraint relations again. First, we have $\prod_{v,p}A(v)B(p)F(v,p)=1$ because this is the same as multiplying all star and plaquette operators before adding the DS patch. Second, we have $\left(\prod_{p\in DS}B^{2}(p)\right)\left(\prod_{p\notin DS}B(p)\right)^{2}=1$. Here, the product is over the new plaquettes, which live inside or adjacent to the right or top of the DS patch, with the square of the rest of the old plaquettes that were not removed. Thus, $l^{\prime}=2$ and $r_{l}=2/2=1$. The crucial observation is that the first constraint will remove an operator with order four, for example, one $F(v,p)$, while the second constraint means we are double-counting an operator of order $2$, taken here as $B^{2}(p)$.
@@
Example VI.7 ($\mathbb{Z}_{4}$ on a Torus with NN DSDS patches).
Starting with example VI.6 with one DSDS patch, we can add one more patch as shown in Fig. 44.
@@
Table 7: Table summarizing changes from $\mathbb{Z}_{4}$ code to $\mathbb{Z}_{4}$ and DS half and half contractible code.
@@
Example VI.8 (Torus with half $\mathbb{Z}_{4}$ and half $DS$ phase).
Let us start with a $\mathbb{Z}_{4}$ on a torus. We want to calculate the degeneracy of a phase where half of the torus is a $\mathbb{Z}_{4}$ QD, and the other half is in the Doubled Semion (DS) phase. Since here the DS lives on a cylinder, the boundary conditions will be different from just a DS patch.
@@
(b) Minimal lattice model for the configuration on the right. Compare with Example VI.11, which uses a different method.
@@
This can be understood intuitively since the full $\mathbb{Z}_{4}$ loops can no longer traverse the torus as some of them get identified or confined when they pass through the DS phase. Another way to calculate the $\operatorname{GSD}$ will make this clearer in VI.2. We saw in example VI.6 that when the DS phase occupied a contractible region, the degeneracy was the same as without the DS phase. The stabilizer counting using minimal models is then a powerful tool for exploring modifications to topological orders in a sequential way. Paired with Alg. 2, one can imagine using a computer program to automate the construction of topological quantum error-correction codes.
VI.2 GSD using Macroscopic Pants Decomposition
Dimensional counting is useful for deriving the logical properties, such as ground state degeneracy, from microscopic configurations, as we saw in the last subsection. This subsection offers a complementary, top-down rather than bottom-up, and intuitive methodology, involving only high-level macroscopic information, for calculating the $\operatorname{GSD}$ by counting the number of non-contractible logical loops of anyons in our model. This is based on Ref. 120. We use these methods to explore and validate results about the introduced composite codes.
Example VI.9 ($D(\mathbb{Z}_{2})$ on a Cylinder with two smooth boundaries).
Let’s take the cylinder with both smooth boundary conditions, as shown in Fig. 46. While there are two non-contractible loops of anyons around the circumference of the cylinder, only the $Z$ loop is non-trivial. Further, an $X$ string between the two smooth boundaries of the cylinder is non-contractible and non-trivial. The $Z$ loop and the $X$ string anticommute, and we have $\operatorname{GSD}=2$. Evidently, all we needed was the information about the topological surface the theory lives on, along with the types of boundaries it has. The specific microscopic details were not used.
@@ VI.2 GSD using Macroscopic Pants Decomposition
In topology, it is important to systematically classify how arbitrary surfaces are constructed in terms of simpler primitive structures. For example, how standard spheres and holes can be glued to form any connected and simply-connected 2-complex is a well-studied problem [6]. In this spirit, this section uses the results of Ref. 120 to validate results in Sec. IV.4.4.
The construction relies on two facts. First, as shown in Fig. 48, any orientable 2D manifold can be decomposed into caps, cylinders, and pants [87, 120]. Second, there is a correspondence between domain walls (boundaries) and a tunneling or transfer matrix that describes the domain wall’s action on anyons [120].
@@
Systematically, the method can be described as follows. First, we begin by taking as an input a 2D surface $\mathcal{M}$ that can have multiple spatial boundaries $\partial\mathcal{M}_{i}$. The surface may also be divided into different topological phases with appropriate domain walls $W^{j}$ between them. Boundaries are described as domain walls with the vacuum $W^{i}$ for uniformity. Second, we decompose the 2D surface using pants decomposition into the 2D components described in Fig. 48. This will isolate different bulk phases that have domain walls between them into different elementary 2D components. Multiple pants decompositions exist for the same configuration, and it is not crucial to use the minimal one. The different decompositions are related by a set of topological moves that will naturally not change the topology of the surface nor affect the $\operatorname{GSD}$ calculated here [87, 120]. We also do not discuss how this decomposition is formally carried out, since in all the examples, finding a decomposition is relatively easy. Third, we attach to every 2D component its appropriate tensor from Fig. 48. The tensors take values in the anyon labels of the bulk (boundary) of their respective topological phases. In this step, we also need to choose labels for the indices. Two indices that are connected should have the same label. In the end, no free indices should remain. Fourth, the $\operatorname{GSD}$ will be the scalar resulting from contracting the indices with the same label. This process is better illustrated with examples.
VI.2.1 Examples using Pants Decomposition
Example VI.10 ($\mathbb{Z}_{2}$ on a Torus with a non-contractible twist II).
@@ VI.2.1 Examples using Pants Decomposition
agreeing with Ref. 144 and our result using stabilizers model in Ex. VI.5.
@@
Example VI.11 ($\mathbb{Z}_{4}$ with a non-contractible DS patch).
@@
which confirms the results of Ex. VI.8.
@@
Example VI.12 ($D(\mathbb{Z}_{4})$ with a one non-contractible DS patch).
@@
All the domain walls between two $\mathbb{Z}_{4}$ phases are trivial and denoted here by $W^{1}$. The domain wall from $\mathbb{Z}_{4}$ to $DS$ is denoted by $W$, which is given in Ex. VI.11. Finally, we attach to the pants the $D(\mathbb{Z}_{4})$ fusion coefficients $N_{i,j}^{k}=\delta_{i+j,k}$ and to the cap the delta operator. The $\operatorname{GSD}$ is then:
@@
where in the second step we used $N_{i,a}^{a}=\delta_{i,1}$. In the last step, we just summed over the anyon types of $D(\mathbb{Z}_{4})$. This agrees with the results of Ex. VI.6.
Example VI.13 ($D(\mathbb{Z}_{4})$ on a torus with a $N$ DS patches).
@@
We note that $W^{\dagger}_{x,1}$ is only non-zero when $x=1$ or $x=e^{2}m^{2}$ as can be seen from Eq. (68). We then have only two choices for the indices $b_{i}$, and there are N1N-1 of them (since $b_{N}$ is dependent). This gives $2^{N-1}$ choices. After picking the free N1N-1 b’s, picking any arbitrary choice for $a_{1}$ from the $16$ anyons of $\mathbb{Z}_{4}$ will determine the rest of $a^{\prime}s$ in addition to $b_{N}$. This gives $16\times 2^{N-1}$ choices. This agrees with Ex. VI.7.
@@
Example VI.14 (DS Patches in a $\mathbb{Z}_{4}$ bulk with smooth boundaries).
@@
In the second line, we enforced the delta functions. In the third line, since $a,c\in\{1,e^{2}m^{2}\}$, the only pure flux $b$ that can appear as (ac)(a-c) is the vacuum $1$. Finally, summing over the two choices of $a$ gives us $\operatorname{GSD}=2$. The case with $N$ patches can be treated similar to Example VI.13
Figure 53: Pants decomposition of the $\mathbb{Z}_{4}$ toric code with two contractible DS patches and a smooth boundary.
@@ VII Conclusion
In this work, we have defined a new family of Abelian quantum error correcting codes. In addition to conventional boundaries of quantum doubles [26, 12], our codes contain finite-sized condensates of a twisted quantum double of the parent quantum double. We have explicitly illustrated this using the group $G=\mathbb{Z}_{4}$ and generalizing the DS global condensation[66] to a spatially local one. We derived the logical operators and showed that, in addition to whatever background logical operators of the parent $D(\mathbb{Z}_{4})$, N1N-1 logical qubits are added via $N$ patches. This result explicitly shows how topological qubits of distinct dimensions (two and four) can co-exist in a single code. We have described why, from the perspective of simplifying and improving error correction in the near term, this approach is desirable.
@@
During the final stages of writing this manuscript, the authors became aware of Ref. [189], which analyzes phase transitions between DS and $\mathbb{Z}_{4}$ phases. While complementary, our work focuses on the quantum error-correcting codes derived from the judicious use of these phases. It will be interesting to investigate the intersection of these two concepts in future studies.